OpenStudy (anonymous):

Find the general antiderivative of the function below. f(t) = (t^4-1)/t

OpenStudy (anonymous):

\[f(t)=\frac{ t^{4} -1 }{ t }\]

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

\[f(t)=t^3-\frac{1}{t}\]

OpenStudy (anonymous):

Okay, so the general derivative would be:\[\frac{ t^{4} }{ 4 }-\ln(t)\]

OpenStudy (anonymous):

+C sorry and absolute value on the ln.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yeah that is it' yw, but you did all the work